What is the arrival time, completion time, burst time, turnaround time, and the waiting time for Multiple-Level Queues Scheduling?
Answer #1

Answer:----------

Multiple-Level Queues Scheduling:------------

Arrival time:----- Time at which the process arrives in the ready queue.

Completion time:---- Time at which process completes its execution.

Burst time:---- Time required by a process for CPU execution.

Turnaround time:---- The interval from the time of submission of a process to the time of completion is the turnaround time.Turnaround time is the sum of the periods spent waiting to get into memory, waiting in the ready queue, executing on the CPU, and doing I/O.
Turnaround Time = Completion Time - Arrival Time

Waiting time :---- The amount of time that a process spends waiting in the ready queue. Waiting time is the sum of the periods spent waiting in the ready queue.
Waiting time = Turnaround time – Burst time
Waiting time = Completion time – Arrival time – Burst Time

Example of Multiple-Level Queues Scheduling:----------
Consider below table of three processes under Multilevel queue scheduling.Queue number denotes the queue of the process.

Process Arrival Time CPU Burst Time Queue Number
P1 0 4 1
P2 0 4 1
P3 8 7 2

Priority of queue 1 is greater than queue 2. queue 1 uses Round Robin (Time Quantum = 2) and queue 2 uses FCFS.

Below is the gantt chart of the problem :

P1 P2 P1 P2 P3

0----------------2-----------------4------------------6-----------------8-----------------15

At starting both queues have process so process in queue 1 (P1, P2) runs first (because of higher priority) in the round robin fashion and completes after 8 units then process in queue 2 (P3) starts running (as there is no process in queue 1) and another 7 units completes its execution.

Completion Time of P1 :-- 6
Completion Time of P2 :-- 8
Completion Time of P3 :-- 15

Turnaround Time of P1 :-- 6 - 0 = 6
Turnaround Time of P2 :-- 8 - 0 = 8
Turnaround Time of P3 :-- 15 - 8 = 7

Waiting Time of P1 :-- 6 - 4 = 2
Waiting Time of P2 :-- 8 - 4 = 4
Waiting Time of P3 :-- 15 - 7 = 8